OpenStudy (anonymous):

2x - 3(x + 4) = 17 x = ?

alones (alones):

2x−3(x+4)=17 Step 1: Simplify both sides of the equation. 2x−3(x+4)=17 −x−12=17 Step 2: Add 12 to both sides. −x−12+12=17+12 −x=29 Step 3: Divide both sides by -1. −x/−1=29/−1 x=−29

OpenStudy (anonymous):

thank you